OpenStudy (anonymous):

Solve tan(x)(tanx + 1) = 0 .

OpenStudy (psymon):

Well, this is already given to you in factored form. When you have everything in factored form, you can take each factor and set them each to 0 independent of one another. So I could say tanx = 0 tanx + 1 = 0 And then solve for x in both.

OpenStudy (anonymous):

the first is πn

OpenStudy (psymon):

That's one of the two answers for the first one, unless you can't include 0.

OpenStudy (anonymous):

the next is x = 1/4 (4πn - π)

OpenStudy (psymon):

Works for me :3

OpenStudy (anonymous):

would that mean its x = πn, and x = π/4 plus or minus nπ?

OpenStudy (anonymous):

wait, am I wrong @Psymon ?

OpenStudy (anonymous):

@cambrige @abb0t was I wrong? x = πn, and x = π/4 plus or minus nπ

OpenStudy (psymon):

Well, all values where tanx= -1. So 3pi/4 + kpi

OpenStudy (psymon):

And then for tanx = 0, itd be 0 + kpi

OpenStudy (psymon):

Can ya see why?

OpenStudy (anonymous):

not really? x = 3π/4 plus or minus 2πn

OpenStudy (psymon):

Alrighty, hang on.

OpenStudy (psymon):

 So those are the 4 values. The period of the tangent function is pi, so you would just add kpi to each of those.